# Partial Information Rate Decomposition

### v1

@@ Abstract
Partial Information Decomposition (PID) is a principled and flexible method to unveil complex high-order interactions in multi-unit network systems. Though being defined exclusively for random variables, PID is ubiquitously applied to multivariate time series taken as realizations of random processes with temporal statistical structure. Here, to overcome the incorrect depiction of high-order effects by PID schemes applied to dynamic networks, we introduce the framework of Partial Information Rate Decomposition (PIRD). PIRD is formalized applying lattice theory to decompose the information shared dynamically between a target random process and a set of source processes, implemented for Gaussian processes through a spectral expansion of information rates, and demonstrated in practice analyzing time series from large-scale climate oscillations.
@@
Even though the PID is specifically defined for random variables, the data sequences over which it is computed are often multivariate time series that can only be regarded as realizations of random processes. In this case, under the implicit underlying assumption that the processes at hand are stationary and memoryless (i.e., composed by independent and identically distributed (i.i.d.) variables), the PID decomposes the information shared instantaneously by the processes. However, the i.i.d. assumption is typically not tested in practice and is often violated in applications of information decomposition where the analyzed data exhibit temporal correlations Kay et al. (2022); Varley et al. (2023a, b). Here, we warn against the use of PID in processes exhibiting evident temporal statistical structure, showing that the presence of temporal correlations has a profound impact on the multivariate information shared at lag zero by multiple random processes (Fig. 1). On the other hand, applications of the PID to time-lagged variables selected ad-hoc from the processes (e.g.,Luppi et al. (2022); Koçillari et al. (2023); Luppi et al. (2024)) are inherently focused on specific portions of the system dynamics, thus revealing the lack of general and comprehensive approaches to information decomposition that can provide a full account of the dynamical nature of random processes. The present work fills this fundamental gap by formalizing the PID framework for networks of stationary random processes, leveraging the use of entropy rates as basic elements of information decomposition.
@@
We start considering the information-theoretic measure quantifying the overall degree of association between two (possibly vector) discrete-time random processes $X=\{X_{t}\}_{t\in\mathbb{Z}}$ and $Y=\{Y_{t}\}_{t\in\mathbb{Z}}$, which is the mutual information rate (MIR) Duncan (1970) defined as the limit
$I_{X;Y}=\lim_{n\to\infty}\frac{1}{n}I(X_{1},\ldots,X_{n};Y_{1},\ldots,Y_{n}),$ (1)
@@
where $I(\cdot;\cdot)$ denotes the mutual information (MI) between two random variables. Then, considering $X$ as composed by $N$ source processes, $X=\{X_{1},\ldots,X_{N}\}$, and denoting $Y$ as target process, we exploit the mathematical lattice structure defined for the PID Williams and Beer (2010) and apply it to represent the set-theoretic intersection of multiple processes, so as to decompose the MIR as
$I_{X;Y}=\sum_{\alpha\in\mathcal{A}}I^{\delta}_{X_{\alpha};Y},$ (2)
@@
where $X_{\alpha}\subset X$ groups the source processes indexed by $\alpha$, and $\mathcal{A}$ is the collection of all subsets of sources such that no subset is a superset of any other (e.g., $\mathcal{A}=\{\{1\}\{2\},\{1\},\{2\},\{12\}\}$ if $N=2$) Williams and Beer (2010). Eq. (2) achieves a so-called partial information rate decomposition (PIRD) whereby the MIR is expanded as the sum of contributions (information rate atoms $I^{\delta}_{\cdot;\cdot}$) identified by the lattice structure. Moreover, the marginal MIR terms involving any individual source process $X_{i}$ are constructed additively by summing the information rate of the atoms positioned at the level $\{i\}$ and downwards in the lattice according to
$I_{X_{i};Y}=\sum_{\beta\preceq\{i\}}I^{\delta}_{X_{\beta};Y},i=1,\ldots,N$ (3)
@@
$I^{\cap}_{X_{\alpha};Y}=\sum_{\beta\preceq\alpha}I^{\delta}_{X_{\beta};Y}.$ (4)
@@
Then, once the redundancy rate is known, the information rate associated to all atoms can be retrieved via Möbius inversion of (4). When $N=2$ source processes are considered, the PIRD atoms identify the redundant, unique and synergistic information rates (respectively, $R_{Y;X}\vcentcolon=I^{\delta}_{X_{\{1\}\{2\}};Y}$, $U_{Y;X_{1}}\vcentcolon=I^{\delta}_{X_{\{1\}};Y}$ and $U_{Y;X_{2}}\vcentcolon=I^{\delta}_{X_{\{2\}};Y}$, and $S_{Y;X}\vcentcolon=I^{\delta}_{X_{\{12\}};Y}$), while for $N>2$ such contributions may result from coarse-graining approaches that sum the information rate of multiple atoms Rosas et al. (2020). Being explicitly defined for random processes, the PIRD defined by (2,4) generalizes previous approaches applying PID to random variables arbitrarily selected from the processes: in the absence of temporal correlations, the PIRD reduces to the static PID decomposing the instantaneous information shared by the processes, $I(X_{t};Y_{t})$; in the case of strictly causal processes with the target not sending information to the sources, it reduces to the PID applied to the joint transfer entropy from all sources to target, $T_{X\rightarrow Y}=I({X_{<t};Y_{t}|Y_{<t}})$ ($X_{<t}=\lim_{k\to\infty}(X_{t-k},\ldots,X_{t-1})$) Sparacino et al. (2025).
@@
The operationalization of the function needed to capture the notion of redundancy rate can follow several different strategies proposed to define redundancy among random variables. These strategies differ depending on the philosophy followed to satisfy the desired properties (e.g., decision-, game-, information-theoretic), on the nature (continuous or discrete) of the analyzed variables, and on assumptions made about their distribution (e.g., Gaussian) Williams and Beer (2010); Barrett (2015); Ince (2017); Gutknecht et al. (2021); Ehrlich et al. (2024). A popular approach is to define redundancy as the ensemble (time-domain) average of a given local redundancy measure obtained elaborating properly computed pointwise MI values Ince (2017); Finn and Lizier (2018). Here we follow a conceptually similar approach, making it specifically tailored to random processes in that (i) it is applied to MIR instead of MI and (ii) it is implemented through a local representation in frequency rather than in time. Formally, we expand the redundancy rate in the frequency domain as
$I^{\cap}_{X_{\alpha};Y}=\frac{1}{2\pi}\int_{-\pi}^{\pi}i^{\cap}_{X_{\alpha};Y}(\omega)\mathrm{d}\omega,$ (5)
@@
where $i^{\cap}_{\cdot;\cdot}$ is the spectral redundancy rate defined at each normalized angular frequency $\omega\in[-\pi,\pi]$ according to the minimum MI principle Barrett (2015) applied to the spectral MIR terms indexed by $\alpha=\{\alpha_{1},\ldots,\alpha_{J}\}$, i.e.
$i^{\cap}_{X_{\alpha};Y}(\omega)=\min_{j=1,\ldots,J}i_{X_{\alpha_{j}};Y}(\omega).$ (6)
In the case of two sources, the spectral redundancy rate corresponds to the individual spectral MIR terms for the unique atoms (i.e., $i^{\cap}_{X_{\{k\}};Y}(\omega)=i_{X_{k};Y}(\omega),k=1,2$), to their minimum for the redundant atom (i.e., $i^{\cap}_{X_{\{1\}\{2\}};Y}(\omega)=\min\{i_{X_{1};Y}(\omega),i_{X_{2};Y}(\omega)\}$), and to the overall spectral MIR for the synergistic atom (i.e., $i^{\cap}_{X_{\{12\}};Y}(\omega)=i_{X_{1},X_{2};Y}(\omega)$). Each spectral MIR is computed from the elements of the power spectral density of the joint process $S=\{X,Y\}=\{X_{1},\cdots X_{N},Y\}$, defined as $P_{S}(\omega)=\mathfrak{F}\{R_{S}(k)\}$, with $R_{S}(k)=\mathbb{E}[S_{n-k}^{\intercal}S_{n}]$ being the covariance of $S$ Sparacino et al. (2024). For instance, if $N=2$ the individual and joint spectral MIR functions are
$i_{X_{k};Y}(\omega)$ $=\frac{1}{2}\log\frac{P_{X_{k}}(\omega)P_{Y}(\omega)}{|P_{[X_{k}Y]}(\omega)|},k=1,2,$ (7)
$i_{X_{1},X_{2};Y}(\omega)$ $=\frac{1}{2}\log\frac{|P_{[X_{1}X_{2}]}(\omega)|P_{Y}(\omega)}{|P_{S}(\omega)|},$
where $P_{[X_{k}Y]}$ and $P_{[X_{1}X_{2}]}$ are sub-matrices of the full spectral density $P_{S}$. Importantly, for Gaussian processes each spectral MIR is connected to its time-domain form by the spectral integration property (e.g., $I_{X_{k};Y}=\int_{-\pi}^{\pi}i_{X_{k};Y}(\omega)\mathrm{d}\omega$) Sparacino et al. (2024). Exploiting this property, in Sparacino et al. (2025) we show that there is a close connection between the spectral partial information and redundancy rates defined over a lattice structure identified at each frequency $\omega$ on the one side, and the time-domain partial information and redundancy rates defined here in (4) on the other side.
@@
As an illustrative example of PIRD, we consider a simple network system whose connections are mapped by the vector random process $S=\{X_{1},X_{2},Y\}$ with dynamics determined by the stochastic equations
$X_{1,n}$ $=a_{1}X_{1,n-1}+b_{1}X_{2,n-1}+U_{1,n},$ (8)
$X_{2,n}$ $=a_{2}X_{2,n-1}+b_{2}X_{1,n-1}+U_{2,n},$
$Y_{n}$ $=c_{1}X_{1,n-1}+c_{2}X_{2,n-1}+W_{n},$
@@
where $W$ and $U_{k}$ ($k=1,2$) are white and stationary innovation processes formed by i.i.d. standard Gaussian variables ($W_{n},U_{k,n}\sim\mathcal{N}(0,1)$) with covariances $R_{U_{1}U_{2}}=\mathbb{E}[U_{1,n}U_{2,n}]$ and $R_{WU_{k}}=\mathbb{E}[W_{n}U_{k,n}]$. As depicted in Fig. 1a, the source processes $X_{1}$ and $X_{2}$ present internal dynamics modulated by the coefficients $a_{1},a_{2}$ (orange arrows), and causal interactions are set between the sources and towards the target $Y$ via the coefficients $b_{1},b_{2}$ and $c_{1},c_{2}$ (green arrows); all time-delayed interactions occur at lag 1, while instantaneous (non-delayed) interactions are determined by the innovation covariances (red dashed arrows).
The system (8) is analyzed by computing both PID and PIRD as a function of a modulation parameter $d\in[0,1]$ that acts on the strength of instantaneous and time-lagged effects. First, we study (8) in the absence of instantaneous effects ($R_{U_{1}U_{2}}=R_{WU_{k}}=0$), letting the strength of causal interactions to increase progressively ($c_{1}=c_{2}=d$), and the strength of internal dynamics to decrease progressively ($a_{1}=a_{2}=0.8(1-d)$), with fixed coupling between the sources ($b_{1}=b_{2}=0.1$). The analysis (Fig. 1b) documents that the target shares information at lag zero with the sources as soon as causal interactions are set ($I(Y_{n};X_{n})>0$ when $d>0$) even if instantaneous effects are kept to zero; this information does not exhibit a monotonic trend, and the PID suggests that it is seen as prevalently redundant ($R(Y_{n};X_{n})>S(Y_{n};X_{n})$). On the other hand, the PIRD captures the increasing rate of information shared by the target and source processes, and documents the emergence of net synergy consequent to the increased strength of the common child relations $X_{1}\rightarrow Y$ and $X_{2}\rightarrow Y$; the unique terms are null in this simulation where the two sources contribute equally to the target. In the second simulation setting, the system is investigated reproducing a transition from purely instantaneous to exclusively time-lagged interactions among the processes, achieved decreasing progressively the strength of zero lag effects ($R_{U_{1}U_{2}}=R_{WU_{2}}=0.25(1-d)$, $R_{WU_{1}}=0.5(1-d)$) while increasing the strength of internal dynamics and causal interactions ($a_{1}=a_{2}=0.2d$, $b_{1}=b_{2}=0.1d$, $c_{1}=c_{2}=0.6d$). The results in Fig. 1c confirm that the PIRD reduces to the PID in the absence of time-lagged effects ($d=0$), but the injection of temporal correlations into the processes makes the two approaches to differ substantially: whilst the information shared between $Y_{n}$ and $\{X_{1,n},X_{2,n}\}$ decreases progressively together with its unique, redundant and synergistic contributions, the overall information rate between $Y$ and $\{X_{1},X_{2}\}$ initially decreases due to diminishing unique contribution of $X_{1}$ and then increases substantially due to the rise of both redundant and synergistic effects; net synergy emerges for $d>0.7$, and the unique terms are both nonzero for $d>0.5$.

### v3

@@ Abstract
Partial Information Decomposition (PID) is a principled and flexible method to unveil complex high-order interactions in multi-unit network systems. Though being defined exclusively for random variables, PID is ubiquitously applied to multivariate time series taken as realizations of random processes with temporal statistical structure. Here, to overcome the incorrect depiction of high-order effects by PID schemes applied to dynamic networks, we introduce the framework of Partial Information Rate Decomposition (PIRD). PIRD is first formalized applying lattice theory to decompose the information shared dynamically between a target random process and a set of source processes, and then implemented for Gaussian processes through a spectral expansion of information rates. The new framework is validated in simulated network systems and demonstrated in the practical analysis of time series from large-scale climate oscillations.
@@
Even though the PID is specifically defined for random variables, the data sequences over which it is computed are often multivariate time series that can only be regarded as realizations of random processes. In this case, under the implicit underlying assumption that the processes at hand are stationary and memoryless (i.e., composed by independent and identically distributed (i.i.d.) variables), the PID decomposes the information shared instantaneously by the processes. However, the i.i.d. assumption is typically not tested in practice and is often violated in applications of information decomposition where the analyzed data exhibit temporal correlations Kay et al. (2022); Varley et al. (2023a, b). Here, we warn against the use of PID in processes exhibiting evident temporal statistical structure, showing that the presence of temporal correlations has a profound impact on the multivariate information shared at lag zero by multiple random processes (Fig. 1). On the other hand, applications of the PID to time-lagged variables selected ad-hoc from the processes (e.g.,Luppi et al. (2022); Koçillari et al. (2023); Luppi et al. (2024)) are inherently focused on specific portions of the system dynamics, thus revealing the lack of general and comprehensive approaches to information decomposition that can provide a full account of the dynamical nature of random processes. The present work fills this fundamental gap first by formalizing the PID framework for networks of random processes through the use entropy rates as basic elements of information decomposition, and then developing a computationally reliable implementation of the approach based on a new definition of redundancy valid for information rates.
@@
We start considering the information-theoretic measure quantifying the overall degree of association between two discrete-time stationary random processes $X=\{X(t_{n})\}_{n\in\mathbb{Z}}$ and $Y=\{Y(t_{n})\}_{n\in\mathbb{Z}}$, which is the mutual information rate (MIR) Duncan (1970) defined as
$I_{X;Y}=\lim_{n\to\infty}\frac{1}{n}I(X(t_{1}),\ldots,X(t_{n});Y(t_{1}),\ldots,Y(t_{n})),$ (1)
@@
where $I(\cdot;\cdot)$ denotes the mutual information (MI) between two random variables. Then, considering X as a vector of $N$ source processes, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}=\{X_{1},\ldots,X_{N}\}$, and denoting $Y$ as the scalar target process, we exploit the mathematical lattice structure defined for the PID Williams and Beer (2010) and apply it to represent the set-theoretic intersection of multiple processes, so as to decompose the MIR as
$I_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}};Y}=\sum_{\alpha\in\mathcal{A}}I^{\delta}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha};Y},$ (2)
@@
where $_{\alpha}\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}$ X groups the source processes indexed by $\alpha$, and $\mathcal{A}$ is the collection of all subsets of sources such that no subset is a superset of any other (e.g., $\mathcal{A}=\{\{1\}\{2\},\{1\},\{2\},\{12\}\}$ if $N=2$) Williams and Beer (2010). Eq. (2) achieves a so-called partial information rate decomposition (PIRD) whereby the MIR is expanded as the sum of contributions (information rate atoms $I^{\delta}_{\cdot;\cdot}$) identified by the lattice structure. Moreover, the marginal MIR terms involving any individual source process $X_{i}$ are constructed additively by summing the information rate of the atoms positioned at the level $\{i\}$ and downwards in the lattice according to
$I_{X_{i};Y}=\sum_{\beta\preceq\{i\}}I^{\delta}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\beta};Y},i=1,\ldots,N$ (3)
@@
$I^{\cap}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha};Y}=\sum_{\beta\preceq\alpha}I^{\delta}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\beta};Y}.$ (4)
@@
Then, once the redundancy rate is known, the information rate associated to all atoms can be retrieved via Möbius inversion of (4). When $N=2$ source processes are considered, the PIRD atoms identify the redundant, unique and synergistic information rates (respectively, $R_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}\vcentcolon=I^{\delta}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\{1\}\{2\}};Y}$, $U_{Y;X_{1}}\vcentcolon=I^{\delta}_{X_{\{1\}};Y}$ and $U_{Y;X_{2}}\vcentcolon=I^{\delta}_{X_{\{2\}};Y}$, and $S_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}\vcentcolon=I^{\delta}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\{12\}};Y}$), while for $N>2$ such contributions may result from coarse-graining approaches that sum the information rate of multiple atoms Rosas et al. (2020). Being explicitly defined for random processes, the PIRD defined by (2,4) generalizes previous approaches applying PID to random variables arbitrarily selected from the processes: in the absence of temporal correlations, the PIRD reduces to the static PID decomposing the instantaneous information shared by the processes, $I(X(t_{n});Y(t_{n}))$; in the case of strictly causal processes with the target not sending information to the sources, it reduces to the PID applied to the joint transfer entropy from all sources to target, $T_{\textbf{X}\rightarrow Y}=I(\textbf{X}(t_{<n});Y(t_{n})|Y(t_{<n}))$, where $\textbf{X}(t_{<n})=\lim_{k\to\infty}(\textbf{X}(t_{n-k}),\ldots,\textbf{X}(t_{n-1})$) Sparacino et al. (2025a).
@@
The operationalization of the function needed to capture the notion of redundancy rate can follow several different strategies proposed to define redundancy among random variables. These strategies differ depending on the philosophy followed to satisfy the desired properties (e.g., decision-, game-, information-theoretic), on the nature (continuous or discrete) of the analyzed variables, and on assumptions made about their distribution (e.g., Gaussian) Williams and Beer (2010); Barrett (2015); Ince (2017); Gutknecht et al. (2021); Ehrlich et al. (2024). A popular approach is to define redundancy as the ensemble (time-domain) average of a given local redundancy measure obtained elaborating properly computed pointwise MI values Ince (2017); Finn and Lizier (2018). Here we follow a conceptually similar approach, making it specifically tailored to random processes in that (i) it is applied to MIR instead of MI and (ii) it is implemented through a local representation in frequency rather than in time. Formally, we define the redundancy rate of the atom $\alpha=\{\alpha_{1},\ldots,\alpha_{J}\}\in\mathcal{A}$ as
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I^{\cap}_{\textbf{X}_{\alpha};Y}:=}\frac{1}{2\pi}\int_{-\pi}^{\pi}i^{\cap}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha};Y}(\omega)\mathrm{d}\omega,$ (5)
@@
where $i^{\cap}_{\cdot;\cdot}(\omega)$ is the pointwise (local) spectral redundancy rate defined at each normalized angular frequency $\omega\in[-\pi,\pi]$ according to the minimum MI principle Barrett (2015) applied to the spectral MIR terms $i_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}};Y}(\omega)$:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i^{\cap}_{\textbf{X}_{\alpha};Y}(\omega):=}\min_{j=1,\ldots,J}i_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}};Y}(\omega).$ (6)
To compute the redundancy rate and solve the PIRD it is necessary to provide a proper spectral MIR function to use in (6). Here, we exploit the linear representation of multivariate Gaussian processes, for which the spectral MIR between any two analyzed processes can be computed from the elements of the power spectral density of the joint process ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}=\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}},Y\}=\{X_{1},\cdots X_{N},Y\}$, defined as the $(N+1)\times(N+1)$ matrix $\textbf{P}_{\textbf{S}}(\omega)=\mathfrak{F}\{\textbf{R}_{\textbf{S}}(k)\}$, with $\textbf{R}_{\textbf{S}}(k)=\mathbb{E}[\textbf{S}^{\intercal}(t_{n-k})\textbf{S}(t_{n})]$ being the covariance of S Sparacino et al. (2025b). Specifically, if S is a jointly stationary Gaussian process, the MIR between the target $Y$ and the source $_{\alpha_{j}}\in\textbf{X}_{\alpha}\subseteq\textbf{X}$ X can be expanded in the frequency domain as $I_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}};Y}=\frac{1}{2\pi}\int_{-\pi}^{\pi}i_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}};Y}(\omega)\mathrm{d}\omega$, where the spectral MIR function is Chicharro (2011)
$i_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}};Y}(\omega)=\frac{1}{2}\log\frac{|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{P}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}}(\omega)|P_{Y}(\omega)}{|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{P}}_{[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}Y]}(\omega)|},$ (7)
being ||$\cdot$|| ⋅ | the matrix determinant and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{P}}_{[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}Y]}$ the sub-matrix of the full spectral density $\textbf{P}_{\textbf{S}}$ relevant to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}$ and $Y$. For instance, in the case of two sources, the spectral redundancy rates obtained via (6) are $i^{\cap}_{X_{\{k\}};Y}(\omega)=i_{X_{k};Y}(\omega)$ for the unique atoms ($k=1,2$), $i^{\cap}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\{1\}\{2\}};Y}(\omega)=\min\{i_{X_{1};Y}(\omega),i_{X_{2};Y}(\omega)\}$ for the redundant atom, and $i^{\cap}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\{12\}};Y}(\omega)=i_{X_{1},X_{2};Y}(\omega)$ for the synergistic atom, where the spectral MIR is computed via (7) respectively with $\alpha=\{k\}$, $\alpha=\{1\}\{2\}$, and $\alpha=\{12\}$.
When implemented by means of the spectral MIR (7), the redundancy rate defined in (5,6) holds only for stationary Gaussian processes admitting a spectral representation Chicharro (2011). Moreover, due to the lack of additivity, the redundancy rate (5) differs from the minimum time-domain MIR; in Sparacino et al. (2025a) we show that $I^{\cap}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha};Y}\leq\min_{j=1,\ldots,J}I_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}};Y}$, denoting that applying the minimum MI principle at the local level in the spectral domain yields more conservative redundancy rates than working at the global level in the time domain. Nevertheless, the proposed redundancy rate preserves non-negativity, satisfies the main axioms of redundancy measures (i.e. symmetry, self-redundancy, and monotonicity) Williams and Beer (2010), and offers the interesting possibility to perform PIRD focusing on predefined frequency bands with practical meaning Sparacino et al. (2025a).
@@
To validate PIRD in an illustrative example, we consider a simple network system whose connections are mapped by the vector random process ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}=\{X_{1},X_{2},Y\}$ with dynamics determined by the stochastic equations
$X_{1}(t_{n})$ $=a_{1}X_{1}(t_{n-1})+b_{1}X_{2}(t_{n-1})+U_{1}(t_{n}),$ (8)
$X_{2}(t_{n})$ $=a_{2}X_{2}(t_{n-1})+b_{2}X_{1}(t_{n-1})+U_{2}(t_{n}),$
$Y(t_{n})$ $=c_{1}X_{1}(t_{n-1})+c_{2}X_{2}(t_{n-1})+W(t_{n}),$
@@
where $W$ and $U_{k}$ are white and stationary innovation processes formed by i.i.d. standard Gaussian variables ($W(t_{n}),U_{k}(t_{n})\sim\mathcal{N}(0,1)$, $k=1,2$) whose covariances $R_{U_{1}U_{2}}=\mathbb{E}[U_{1}(t_{n})U_{2}(t_{n})]$ and $R_{WU_{k}}=\mathbb{E}[W(t_{n})U_{k}(t_{n})]$ determine the strength of instantaneous (non-delayed) interactions (Fig. 1a, red dashed lines). The source processes $X_{1}$ and $X_{2}$ present internal dynamics modulated by the coefficients $a_{1},a_{2}$ (orange arrows), and causal interactions are set between the sources and towards the target $Y$ via the coefficients $b_{1},b_{2}$ and $c_{1},c_{2}$ (green arrows); all these time-delayed interactions occur at lag 1. The balance between the strength of the causal mechanisms directed from sources to target ($X_{1}\rightarrow Y\leftarrow X_{2}$) or exerted between the sources ($X_{1}\leftrightarrow X_{2}$) determines the prevalence of common target effects or common drive effects, which are known to induce respectively dominance of synergy or dominance of redundancy Mijatovic et al. (2025).